# Agent-RLVR: Training Software Engineering Agents via Guidance and Environment Rewards

### v1

@@ 1 Introduction
To address these issues, we introduce Agent-RLVR, a framework for training language model agents that relies solely on rewards produced by the agent-environment interaction. We demonstrate the success of the method in software engineering (SWE) tasks. We choose this setting for several reasons: the setting has objectively verifiable solutions (e.g. passing unit tests), previous work such as SWE-Bench (chowdhury2024swebenchverified, ) has provided standardized evaluation infrastructure, and advances in this area offer significant practical value for improving developer productivity. The core of our method is the incorporation of agent guidance—teacher guidance that provides targeted information about failing code patches, file contexts, or test expectations. These hints steer the agent towards successful trajectory by providing these diverse information clues, similar to how a junior software engineer may receive guidance from a teacher or tech lead when first exploring a new codebase.
@@ 2.2 Extending RLVR to Agentic Tasks with Agent-RLVR
1:$M_{\theta}$ = agent model
2:$T$ = teacher
3:$\mathcal{D}$ = dataset of issue descriptions, repositories, and test suites
@@
5:$\mathcal{D}_{\text{correct}}=\emptyset$, $\mathcal{D}_{\text{incorrect}}=\emptyset$
6:for each $(x_{i},r_{i},t_{i})\in\mathcal{D}$ do
7:     $y_{i}=M_{\theta}(x_{i},r_{i})$
8:     $f_{i}=\text{evaluate}(y_{i},t_{i})$
9:     if $f_{i}=\text{pass}$ then $\mathcal{D}_{\text{correct}}=\mathcal{D}_{\text{correct}}\cup\{(x_{i},r_{i},y_{i})\}$
10:     else $e_{i}=\text{getError}(y_{i},t_{i})$, $\mathcal{D}_{\text{incorrect}}=\mathcal{D}_{\text{incorrect}}\cup\{(x_{i},r_{i},y_{i},e_{i})\}$
@@
14:$\mathcal{D}_{\text{guided}}=\emptyset$
15:for each $(x_{i},r_{i},y_{i},e_{i})\in\mathcal{D}_{\text{incorrect}}$ do
16:     $g_{i}=T(x_{i},r_{i},y_{i},e_{i})$
17:     $\mathcal{D}_{\text{guided}}=\mathcal{D}_{\text{guided}}\cup\{(x_{i},r_{i},g_{i},y_{i},e_{i})\}$
18:end for
19:$\mathcal{D}_{\text{pref}}=\emptyset$, $\mathcal{D}_{\text{SFT}}\subset\mathcal{D}$
20:for each $(x_{i},r_{i},g_{i},y_{i},e_{i})\in\mathcal{D}_{\text{guided}}$ do
21:     $y_{i}^{\prime}=M_{\theta}(x_{i},r_{i},g_{i})$
22:     $f_{i}^{\prime}=\text{evaluate}(y_{i}^{\prime},t_{i})$
23:     if $f_{i}^{\prime}=\text{pass}$ then $\mathcal{D}_{\text{pref}}=\mathcal{D}_{\text{pref}}\cup\{(x_{i},r_{i},g_{i},y_{i}^{\prime},y_{i})\}$
24:     end if
25:end for
@@
27:$\mathcal{L}_{\text{SFT}}(\theta)=-\frac{1}{|\mathcal{D}_{\text{SFT}}|}\sum_{(x_{i},r_{i},y_{i}^{*})\in\mathcal{D}_{\text{SFT}}}\log p_{\theta}(y_{i}^{*}|x_{i},r_{i})$
28:$\theta_{\text{SFT}}=\arg\min_{\theta}\mathcal{L}_{\text{SFT}}(\theta)$
29:$\pi_{\text{ref}}=M_{\theta_{\text{SFT}}}$
30:$\mathcal{L}_{\text{DPO}}(\theta)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{\text{pref}}}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right)\right]$
31:$\theta_{\text{RLVR}}=\arg\min_{\theta}\mathcal{L}_{\text{DPO}}(\theta)$
32:return $M_{\theta_{\text{RLVR}}}$

### v2

@@ 1 Introduction
To address these issues, we introduce Agent-RLVR, a framework for training language model agents that relies solely on rewards produced by the agent-environment interaction. We demonstrate the success of the method in software engineering (SWE) tasks. We choose this setting for several reasons: the setting has objectively verifiable solutions (e.g. passing unit tests), previous work such as SWE-Bench (chowdhury2024swebenchverified, ) has provided standardized evaluation infrastructure, and advances in this area offer significant practical value for improving developer productivity. The core of our method is the incorporation of agent guidance—a multi-agent framework where teacher guidance, which provides targeted information about failing code patches or test expectations, is incorporated during RL training. These hints steer the agent towards successful trajectory by providing these diverse information clues, similar to how a junior software engineer may receive guidance from a teacher or tech lead when first exploring a new codebase.
@@ 2.2 Extending RLVR to Agentic Tasks with Agent-RLVR
1:Input: $M_{\theta}$ = agent model, $T$ = teacher, $\mathcal{D}$ = dataset of issue descriptions, repositories, and test suites, $\mathcal{T}_{correct},\mathcal{T}_{incorrect}=\emptyset$: correct and incorrect agent trajectories
2:Initialize $\mathcal{D}_{RLVR}=\emptyset$
@@
4:for $d_{i}$ in dataset $\mathcal{D}$ do
5:     Generate agent trajectory $t_{i}$: $t_{i}=M_{\theta}(d_{i})$
6:     Generate reward $r_{i}$: $r_{i}=\text{evaluate}(d_{i},t_{i})$
7:     if $r_{i}=1$ then
8:         $(d_{i},t_{i})$rightarrow $\mathcal{T}_{correct}$
9:     else
10:         $(d_{i},t_{i})$rightarrow $\mathcal{T}_{incorrect}$
@@
13:Add to $\mathcal{D}_{RLVR}$: $\{(d_{i},t_{i,j}^{+},t_{i,k}^{-}):r_{i,j}>r_{i,k}\}$triangleright Generate preference pairs from multiple rollouts
@@
15:for $(d_{i},t_{i})$ in $\mathcal{T}_{incorrect}$ do
16:     Generate teacher guidance $g_{i}$: $g_{i}=T(d_{i},t_{i})$
17:     Generate new solution with guidance: $t_{i}^{\prime}=M_{\theta}(d_{i},g_{i})$
18:     Evaluate new solution: $r_{i}^{\prime}=\text{evaluate}(d_{i},t_{i}^{\prime})$
19:     if $r_{i}^{\prime}=1$ then
20:         Add $(d_{i},t_{i}^{\prime},t_{i})$ to $\mathcal{D}_{RLVR}$
21:     end if
22:end for
@@
24:$\mathcal{D}_{SFT}\subset\mathcal{D_{RLVR}}$
25:$\mathcal{L}_{\text{SFT}}(\theta)=-\frac{1}{|\mathcal{D}_{\text{SFT}}|}\sum_{(x_{i},r_{i},y_{i}^{*})\in\mathcal{D}_{\text{SFT}}}\log p_{\theta}(y_{i}^{*}|x_{i},r_{i})$
26:$\theta_{\text{SFT}}=\arg\min_{\theta}\mathcal{L}_{\text{SFT}}(\theta)$
27:$\pi_{\text{ref}}=M_{\theta_{\text{SFT}}}$
28:$\mathcal{L}_{\text{DPO}}(\theta)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{RLVR}}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right)\right]$
29:$\theta_{\text{RLVR}}=\arg\min_{\theta}\mathcal{L}_{\text{DPO}}(\theta)$
30:return $M_{\theta_{\text{RLVR}}}$